how do you multiply y?
for example:
5y * y?

Answer by mananth(16949)

In multiplication of terms you multiply the co-efficients and
use the law of indices .( add up the power of the variables)
2y*2y = 2*2*y*y = 4y^2
...
3x*6x^2 = 3*6*x*x*x = 18x^3
...
2a^2*3b= 2*3*a*a*b = 6a^2b
...
m.ananth@hotmail.ca